Montserrat Volcano Observatory

Daily Report
Report for the period 6 pm 3 April
to 6 pm 4 April 1998

The level of volcanic activity over the past 24 hours has remained relatively low.

Between 16:00 yesterday and 16:00 today there were 11 rockfalls and 4 volcano-tectonic earthquakes.

A long GPS occupation was carried out overnight at seven sites. Results were processed today and show that the line between Harris and Hermitage continues to shorten. In other words the Hermitage site is still moving towards the north east.

This morning, a team from the University of Puerto Rico in collaboration with MVO prepared a site at Hermitage for the installation of a gravimeter.

Preparations for the installation of some new seismic stations continued today, mainly in the Silver Hills where a receiver was erected.

A team also went to the Galways area this afternoon to study the deposits of the Boxing Day event and to collect samples.

Dust levels at all monitoring stations were low today.

Residents are reminded to keep listening to Radio Montserrat for any further information.

Montserrat Volcano Observatory